OpenStudy (anonymous):

integral x*sqrt(x^2+2x+10)

OpenStudy (anonymous):

\[\int\limits x.\sqrt(x^2+2x+10)\]

OpenStudy (michele_laino):

please try this change of variable: \[x ^{2}+2x+10=y\] where y is your new variable of integration

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

why?

OpenStudy (mathmath333):

to make it simpler

OpenStudy (anonymous):

oh, you mean substitution?

OpenStudy (michele_laino):

I think that the above substitution doesn't work well!

OpenStudy (anonymous):

\[\sqrt(x^2+2x+1+9) = (x+1)^2+9\]

OpenStudy (anonymous):

i then substitute

OpenStudy (anonymous):

\[t = x+1 <=> t-1 = x\]

OpenStudy (anonymous):

and dx=dt

hartnn (hartnn):

that will be equivalent to integration by parts :) if you're comfortable, then you can directly go to by parts method

OpenStudy (anonymous):

thats where i get stuck

OpenStudy (ksaimouli):

\[\int\limits_{}^{} u*dv= u*v-\int\limits_{}^{}v*du\]

OpenStudy (anonymous):

yea, i tried

OpenStudy (michele_laino):

please, since I can write this identity,: \[x ^{2}+2x+10=(x+2)^{2}+6\] please try this substitution: y=x+2 @Rowa

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

won't that be the same thing i did?

OpenStudy (michele_laino):

oops. I have made an error,: \[x ^{2}+2x+10=(x+1)^{2}+9\] so y=x+1

OpenStudy (anonymous):

yes, i've gotten to that part the integration by parts is complicated for me

OpenStudy (michele_laino):

you should get this: \[\int\limits (y-1)\sqrt{y ^{2}+9} dy\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now integration by parts, which is quite complicated

OpenStudy (anonymous):

i've tried following wolframalpha, but it made things way to complicated

OpenStudy (michele_laino):

so now you can decompose that integral, namely: \[\int\limits y \sqrt{y ^{2}+9}dy-\int\limits \sqrt{y ^{2}+9}dy\]

OpenStudy (anonymous):

ah yes, true

OpenStudy (michele_laino):

ok! Now you can apply this substitution for the first integral: \[y ^{2}+9=z\]

OpenStudy (anonymous):

again? ok

OpenStudy (anonymous):

works well for the left integral, the right one...not so much

OpenStudy (michele_laino):

you should get this: \[\frac{ (x ^{2}+2x+10)^{3/2} }{ 3 }\]

OpenStudy (anonymous):

as answer?

OpenStudy (anonymous):

as final answer?

OpenStudy (michele_laino):

for first integral only!

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thats correct

OpenStudy (michele_laino):

for second integral, try this substitution, please: \[\sqrt{y ^{2}+9}=y+z\] where z is your new variable please try to square both sides of that euation,

OpenStudy (michele_laino):

oops.. of that equation...

OpenStudy (michele_laino):

you should get this: \[y=\frac{ 1 }{ 2 }\frac{ 9-z ^{2} }{ z }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i'm sorry, but i will have to pauze this and continue later

OpenStudy (michele_laino):

so now try to differentiate that equation in order to get dy, please

OpenStudy (anonymous):

Thank you very much, i will continue to apply your instructions when i get back

OpenStudy (michele_laino):

ok!

ganeshie8 (ganeshie8):

thats a pretty neat substitution \(\sqrt{x^2+bx+c} = x+t\)

ganeshie8 (ganeshie8):

\[\int x \sqrt{x^2+2x+10} ~dx\] \(\sqrt{x^2+2x+10} = x+t\) squaring we get \(2x+10 = 2xt + t^2\) \(x = \dfrac{t^2-10}{ 2-2t}\) the integral becomes \[\int \frac{t^2}{8}+\frac{9}{2 (t-1)}-\frac{81}{8 (t-1)^2}+\frac{81}{4 (t-1)^3}-\frac{729}{8 (t-1)^4}+1~dt\] which is easy to deal with

OpenStudy (anonymous):

@Michele_Laino stuck at the second integral

OpenStudy (anonymous):

so\[\sqrt{y ^{2}+9}=y+z\]?

OpenStudy (michele_laino):

yes! Please square both sides, and then solve for y

OpenStudy (michele_laino):

you should get this: \[y=\frac{ 1 }{ 2 }\frac{ 9-z ^{2} }{ z }\]

OpenStudy (anonymous):

how did you come up with \[\sqrt{y ^{2}+9}=y+z\]

OpenStudy (michele_laino):

I write all your steps:

OpenStudy (michele_laino):

\[y ^{2}+9=(y+z)^{2}\] so: \[y ^{2}+9=y ^{2}+2yz+z ^{2}\] then: \[2yz=9-z ^{2}\] finally: \[y=\frac{ 1 }{ 2 }\frac{ 9-z ^{2} }{ z }\]

OpenStudy (anonymous):

ok thanks, but why \[\sqrt{y ^{2}+9}=y+z\]

OpenStudy (anonymous):

i get the first part, the y+z is difficult to understand

OpenStudy (michele_laino):

it is technique of integration!

OpenStudy (anonymous):

oh ok, i will have to learn it

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now, please differentioate that formula above, in order to get dy, namely: \[dy=...dz\]

OpenStudy (michele_laino):

you should get this: \[dy=-\frac{ 1 }{ 2 }\frac{ z ^{2}+9 }{ z ^{2} }dz\]

OpenStudy (anonymous):

thats quick, ok

OpenStudy (michele_laino):

ok! now, please substitute the expressions which we have found, namely y, dy and sqrt(y^2+9) as function of z, into your integral

OpenStudy (michele_laino):

you should get this: \[-\frac{ 1 }{ 4 }\int\limits \frac{ (9+z ^{2})^{2} }{ z ^{3} }dz\]

OpenStudy (anonymous):

what is sqrt((y^2+9) equal to, i mean substituted as

OpenStudy (michele_laino):

oops...note that: \[\sqrt{y ^{2}+9}=y+z=\frac{ 9-z ^{2} }{ 2z }+z=\frac{ 9+z ^{2} }{ 2z }\]

OpenStudy (anonymous):

i think i'm lost

OpenStudy (michele_laino):

where?

OpenStudy (anonymous):

when you used sqrt(t^2+9) = t+z

OpenStudy (anonymous):

i thought you were substituting at first..

OpenStudy (michele_laino):

it is the same, I do that substitution now

OpenStudy (anonymous):

then what is y equal to, if you want to substitute it

OpenStudy (michele_laino):